# TrIM, Triangular Input Movement Systolic Array for Convolutional Neural Networks:

### v2

@@ III Triangular Input Movement Systolic Array
Figure 5: The TrIM architecture and dataflow considering a $5\times 5$ ifmap and a $3\times 3$ kernel: (a) the SA consists of $3\times 3$ PEs and 2 SRBs to enable the diagonal movement of inputs. An extra adder tree, consisting of two adders, accumulates the psums coming from the bottom PEs; (b) the evolution over the different cycles, reporting the computations performed by each PE, the data currently stored into the SRBs and the current output provided by the adder tree. Cells in blue refer to inputs fetched externally, while cells in red indicate inputs reused through a right-to-left movement. Conversely, cells in orange refer to diagonal reuse, either from a SRB or a different PE. Finally, the green boxes refer to the outputs provided by the adder tree. Bold numbers refer to active multiplications. Xs refer to don’t care cases.
@@ III-A The Dataflow
The TrIM-based SA consists of $K\times K$ PEs, where weights are retained at the PE level, while inputs and psums are moved across the array to maximize data reuse. Before any computation, weights are read from the memory and provided to the PEs placed at the top side of the array; then, these are moved from top to down to allow the other weights to be accommodated. Inputs supply the PEs through a novel data movement that can be summarized in three steps: (i) first, inputs are fetched from the main memory and delivered to the PEs; (ii) then, such inputs move from right to left, until they reach the left edge of the array; (iii) finally, they move diagonally towards the upper PEs. According to the width of the ifmap under processing, the leftmost PEs may be connected to Shift Register Buffers (SRBs) with depth $W_{I}-K-1$, which ensure the correct execution of the diagonal movement over time. Steps (i) to (iii) compose a right-triangular shape, hence the name Triangular Input Movement (TrIM). Fig. 4 schematizes a generic architecture coping with the proposed dataflow. Each PE is labelled as $PE_{i,j}$, with $0\leq i,j<K$, and consists of a MAC unit, four registers, and two multiplexers to establish whether the current input is reused from a different PE or not. If reused, inputs can move from right to left (thus from each $PE_{i,j+1}$ to $PE_{i,j}$), or they can be forwarded diagonally. In the latter case, they are first provided to the SRBs from each $PE_{i,0}$, then shifted along such buffers, and finally the inputs stored into the last $K$ registers are supplied to the top $PE_{i-1,j}$ elements. In general, when the generic input is grabbed from the memory, the data $I_{ext}$ is supplied to the PE; when the input stored in the right-adjacent PEs is moved into the current PE, thus the data $I_{R}$ is the target input; finally, when a diagonal movement of the input (either from a bottom PE or from a bottom SRB) has to feed the PE, thus the data $I_{D}$ is provided. Psums are accumulated vertically, thus from each $PE_{i,j}$ to each $PE_{i+1,j}$. Eventually, an adder tree accumulates the psums coming from the the bottom $PE_{K-1,j}$ elements. For very small ifmaps, with $W_{I}\leq 2K$, the number of registers may be lower than $K$, thus the diagonal connections may also interest some of the PEs.
@@
To better understand how TrIM works, we consider the case reported in Fig. 5, where a $5\times 5$ ifmaps, with $I=1,2,...,25$ being the inputs, and a $3\times 3$ kernel, with $W=A,B,...,I$ being the weights, are considered. The equivalent TrIM architecture is made of 3 rows (named $Row_{0}$, $Row_{1}$, and $Row_{2}$), each having 3 PEs. The leftmost PEs belonging to $Row_{1}$ and $Row_{2}$ are connected to SRBs with depth 1. A final adder tree, consisting of two adders, manages the psums coming from the PEs of $Row_{2}$. In what follows, the detailed functionality of the dataflow is explained cycle-by-cycle. Preliminarily, $K$ cycles are needed to load the $K\times K$ weights, arranged in rows of $K$ weights per cycle. After this phase, the actual computations and input movements can start:
  • Cycle 1: $I=1,2,3$, supplied vertically to $Row_{0}$, are multiplied by the weights $W=A,B,C$.
  • Cycle 2: $Row_{0}$ multiplies $I=2,3,4$ with $W=A,B,C$, where $I=2,3$ are reused through right-to-left movements, while $I=4$ is supplied externally. $Row_{1}$ multiplies the new inputs $I=6,7,8$ with $W=D,E,F$, and accumulate these with the psums coming from $Row_{0}$.
  • Cycle 3: $Row_{0}$ multiplies $I=3,4,5$ with $W=A,B,C$, where $I=3,4$ are reused through right-to-left movements, while $I=5$ is supplied externally. $Row_{1}$ multiplies the inputs $I=7,8,9$ with $W=D,E,F$, and accumulate these with the psums coming from $Row_{0}$. While $I=7,8$ are reused through right-to-left movements, $I=9$ is supplied externally. $Row_{2}$ multiplies the new inputs $I=11,12,13$ with $W=G,H,I$, and accumulate these with the psums coming from $Row_{1}$. At this cycle, $SRB_{0}$ receives $I=6$ to be reused later.
  • Cycle 4: $Row_{0}$ multiplies $I=6,7,8$ with $W=A,B,C$, where $I=6$ is reused diagonally from $SRB_{0}$, while $I=7,8$ are reused diagonally from $PE_{1,0}$ and $PE_{1,1}$, respectively. $Row_{1}$ multiplies the inputs $I=8,9,10$ with $W=D,E,F$, and accumulate these with the psums coming from $Row_{0}$. While $I=8,9$ are reused through right-to-left movements, $I=10$ is supplied externally. $Row_{2}$ multiplies the inputs $I=12,13,14$ with $W=G,H,I$, and accumulate these with the psums coming from $Row_{1}$. While $I=12,13$ are reused through right-to-left movements, $I=14$ is supplied externally. At this cycle, $SRB_{0}$ and $SRB_{1}$ receive $I=7$ and $I=11$, respectively. Finally, the psums coming from $PE_{2,0}$, $PE_{2,1}$, and $PE_{2,2}$ are accumulated through the adder tree to get the very first output activation $O_{0,0}$.
  • Cycle 5: $Row_{0}$ multiplies $I=7,8,9$ with $W=A,B,C$, where $I=7,8$ are reused through right-to-left movements, while $I=9$ is reused diagonally from $PE_{1,1}$. $Row_{1}$ multiplies the inputs $I=11,12,13$ with $W=D,E,F$, and accumulate these with the psums coming from $Row_{0}$. $I=11$ is reused diagonally from $SRB_{0}$, while $I=12,13$ are reused diagonally from $PE_{1,0}$ and $PE_{1,1}$, respectively. $Row_{2}$ multiplies the inputs $I=13,14,15$ with $W=G,H,I$, and accumulate these with the psums coming from $Row_{1}$. While $I=13,14$ are reused through right-to-left movements, $I=15$ is supplied externally. At this cycle, $SRB_{0}$ and $SRB_{1}$ receive $I=8$ and $I=12$, respectively. Finally, the psums coming from $PE_{2,0}$, $PE_{2,1}$, and $PE_{2,2}$ are accumulated through the adder tree to get the second output activation $O_{0,1}$.
  • Cycle 6: $Row_{0}$ multiplies $I=8,9,10$ with $W=A,B,C$, where $I=8,9$ are reused through right-to-left movements, while $I=10$ is supplied externally. $Row_{1}$ multiplies the inputs $I=12,13,14$ with $W=D,E,F$, and accumulate these with the psums coming from $Row_{0}$. While $I=12,13$ are reused through right-to-left movements, $I=14$ is reused diagonally from $PE_{2,1}$. $Row_{2}$ multiplies the new inputs $I=16,17,18$ with $W=G,H,I$, and accumulate these with the psums coming from $Row_{1}$. At this cycle, $SRB_{0}$ and $SRB_{1}$ receive $I=11$ and $I=13$, respectively. Finally, the psums coming from $PE_{2,0}$, $PE_{2,1}$, and $PE_{2,2}$ are accumulated through the adder tree to get the third output activation $O_{0,2}$.
  • Cycle 7: $Row_{0}$ multiplies $I=11,12,13$ with $W=A,B,C$, where $I=11$ is reused diagonally from $SRB_{0}$, while $I=12,13$ are reused diagonally from $PE_{1,0}$ and $PE_{1,1}$, respectively. $Row_{1}$ multiplies the inputs $I=13,14,15$ with $W=D,E,F$, and accumulate these with the psums coming from $Row_{0}$. While $I=13,14$ are reused through right-to-left movements, $I=15$ is supplied externally. $Row_{2}$ multiplies the inputs $I=17,18,19$ with $W=G,H,I$, and accumulate these with the psums coming from $Row_{1}$. While $I=17,18$ are reused through right-to-left movements, $I=19$ is supplied externally. At this cycle, $SRB_{0}$ and $SRB_{1}$ receive $I=12$ and $I=16$, respectively. Finally, the psums coming from $PE_{2,0}$, $PE_{2,1}$, and $PE_{2,2}$ are accumulated through the adder tree to get the fourth output activation $O_{1,0}$.
  • Cycle 8: $Row_{0}$ multiplies $I=12,13,14$ with $W=A,B,C$, where $I=12,13$ are reused through right-to-left movements, while $I=14$ is reused diagonally from $PE_{1,1}$. $Row_{1}$ multiplies the inputs $I=16,17,18$ with $W=D,E,F$, and accumulate these with the psums coming from $Row_{0}$. $I=16$ is reused diagonally from $SRB_{0}$, while $I=17,18$ are reused diagonally from $PE_{1,0}$ and $PE_{1,1}$, respectively. $Row_{2}$ multiplies the inputs $I=18,19,20$ with $W=G,H,I$, and accumulate these with the psums coming from $Row_{1}$. While $I=18,19$ are reused through right-to-left movements, $I=20$ is supplied externally. At this cycle, $SRB_{0}$ and $SRB_{1}$ receive $I=13$ and $I=17$, respectively. Finally, the psums coming from $PE_{2,0}$, $PE_{2,1}$, and $PE_{2,2}$ are accumulated through the adder tree to get the fifth output activation $O_{1,1}$.
  • Cycle 9: $Row_{0}$ multiplies $I=13,14,15$ with $W=A,B,C$, where $I=13,14$ are reused through right-to-left movements, while $I=15$ is supplied externally. $Row_{1}$ multiplies the inputs $I=17,18,19$ with $W=D,E,F$, and accumulate these with the psums coming from $Row_{0}$. While $I=17,18$ are reused through right-to-left movements, $I=19$ is reused diagonally from $PE_{2,1}$. $Row_{2}$ multiplies the new inputs $I=21,22,23$ with $W=G,H,I$, and accumulate these with the psums coming from $Row_{1}$. At this cycle, $SRB_{0}$ and $SRB_{1}$ receive $I=16$ and $I=18$, respectively. Finally, the psums coming from $PE_{2,0}$, $PE_{2,1}$, and $PE_{2,2}$ are accumulated through the adder tree to get the sixth output activation $O_{1,2}$.
  • Cycle 10: $Row_{0}$ has completed the computations related to the current ifmap, so it starts to compute a sliding windows from a different ifmap. $Row_{1}$ multiplies the inputs $I=18,19,20$ with $W=D,E,F$, and accumulate these with the psums coming from $Row_{0}$. While $I=18,19$ are reused through right-to-left movements, $I=20$ is supplied externally. $Row_{2}$ multiplies the inputs $I=22,23,24$ with $W=G,H,I$, and accumulate these with the psums coming from $Row_{1}$. While $I=22,23$ are reused through right-to-left movements, $I=24$ is supplied externally. From this cycle, $SRB_{0}$ and $SRB_{1}$ do not need to receive data anymore. Finally, the psums coming from $PE_{2,0}$, $PE_{2,1}$, and $PE_{2,2}$ are accumulated through the adder tree to get the seventh output activation $O_{2,0}$.
  • Cycle 11: $Row_{0}$ and $Row_{2}$ have completed the computations related to the current ifmap, so they work on a different ifmap. $Row_{2}$ multiplies the inputs $I=23,24,25$ with $W=G,H,I$, and accumulate these with the psums coming from $Row_{1}$. While $I=23,24$ are reused through right-to-left movements, $I=25$ is supplied externally. Finally, the psums coming from $PE_{2,0}$, $PE_{2,1}$, and $PE_{2,2}$ are accumulated through the adder tree to get the eighth output activation $O_{2,1}$.
  • Cycle 12: All the rows have completed the computations related to the current ifmap, so they work on a different ifmap. The adder tree provides the final output activation $O_{2,2}$.
@@
The example just presented allows to appreciate how TrIM maximizes inputs utilization. In fact, the number of total memory accesses from the main memory is 29, of which only 4 accesses refer to inputs read more than once. To better spotlight this, let consider the case $I=13$. This input is read once from the main memory (cycle 3) and reused 8 times (from cycle 4 to cycle 9) through right-to-left and diagonal movements.
III-B Differences with WS and RS Dataflows
@@
  • Data redundancy: WS has data redundancy at the main memory level which adds a significant energy overhead. RS has data redundancy at the array level through memory blocks adding still significant energy overhead. While RS moves the redundancy penalty from the off-chip to the on-chip level, TrIM overcomes the input redundancy problem and moreover maximizes the data utilization by recycling the same input as much as possible.
@@
  • Array sizes: WS and TrIM make use of $K\times K$ PEs. On the contrary, RS is a rectangular array with sizes $K\times H_{O}$, thus exhibiting more limitations regarding the design scalability and area efficiency. In addition, the presence of scratch pads at the PE level makes the RS even bigger.
@@ IV An Analytical Model for Systolic Arrays
In order to highlight the advantages provided by TrIM, we built an analytical model to explore and evaluate the design space of TrIM, searching for optimum design points that lead the physical implementation phase. The aim of this model is twofold: (i) providing insights about memory accesses, throughput and local buffers/registers; (ii) comparing TrIM to WS and RS dataflows. Without loss of generality, all the metrics refer to a convolution between an $H_{I}\times W_{I}$ ifmap and a $K\times K$ kernel. However, these can be easily scaled to manage $M$ ifmaps and $N$ filters of $M$ kernels.
@@ IV-B Modelling TrIM Dataflow
$OV=\begin{cases}(W_{I}-K-1)\times(K-1)\times(H_{I}-K)&W_{I}<2K\\ (K-1)^{2}\times(H_{I}-K)&W_{I}\geq 2K\\ \end{cases}$ (14)
@@
The component $4\times K^{2}$ relates to the registers available at the PE level (4 for each PE). The second block $(K-1)\times(W_{I}-K-1)$ refers to the registers that model the SRBs. Finally the extra register equips the adder tree.

### v3

@@ III Triangular Input Movement Systolic Array
Figure 5: The TrIM architecture and dataflow considering a $\times 5$ 5 ifmap and a $\times 3$ 3 kernel: (a) the SA consists of $\times 3$ 3 PEs and 2 SRBs to enable the diagonal movement of inputs. An extra adder tree, consisting of two adders, accumulates the psums coming from the bottom PEs; (b) the evolution over the different cycles, reporting the computations performed by each PE, the data currently stored into the SRBs and the current output provided by the adder tree. Cells in blue refer to inputs fetched externally, while cells in red indicate inputs reused through a right-to-left movement. Conversely, cells in orange refer to diagonal reuse, either from a SRB or a different PE. Finally, the green boxes refer to the outputs provided by the adder tree. Bold numbers refer to active multiplications. Xs refer to don’t care cases. Weights are preloaded through vertical links (not shown in this figure) and keep stationary during all the convolutional computations.
@@ III-A The Dataflow
The TrIM-based SA consists of $K\times K$ PEs, where weights are retained at the PE level, while inputs and psums are moved across the array to maximize data reuse.
Before starting the convolution between the current ifmap and the respective kernel, weights are supplied to the SA. These weights are loaded from top to bottom: in the first cycle, the first group of $K$ weights is supplied to the top row of the array; in the following cycle, the previous group of weights is shifted to the second row, while the first row receives a new group of $K$ weights; and so on, until the entire SA is filled with the $K\times K$ weight kernel. Thus, weights are kept stationary until the completion of all the convolutions.
@@
Inputs supply the PEs through a novel data movement that can be summarized in three steps: (i) first, inputs are fetched from the main memory and delivered to the PEs; (ii) then, such inputs move from right to left, until they reach the left edge of the array; (iii) finally, they move diagonally towards the upper PEs. According to the width of the ifmap under processing, the leftmost PEs may be connected to Shift Register Buffers (SRBs) with depth $W_{I}-K-1$, which ensure the correct execution of the diagonal movement over time. Steps (i) to (iii) compose a right-triangular shape, hence the name Triangular Input Movement (TrIM). Fig. 4 schematizes a generic architecture coping with the proposed dataflow. Each PE is labeled as $PE_{i,j}$, with $0\leq i,j<K$, and consists of a MAC unit, four registers, and two multiplexers to establish whether the current input is reused from a different PE or not. If reused, inputs can move from right to left (thus from each $PE_{i,j+1}$ to $PE_{i,j}$), or they can be forwarded diagonally. In the latter case, they are first provided to the SRBs from each $PE_{i,0}$, then shifted along such buffers, and finally the inputs stored into the last $K$ registers are supplied to the top $PE_{i-1,j}$ elements. In general, when the generic input is grabbed from the memory, the data $I_{ext}$ is supplied to the PE; when the input stored in the right-adjacent PEs is moved into the current PE, thus the data $I_{R}$ is the target input; finally, when a diagonal movement of the input (either from a bottom PE or from a bottom SRB) has to feed the PE, thus the data $I_{D}$ is provided. Algorithm 1 pitches the orchestration of the three movements over time.
Algorithm 1 TrIM Dataflow
[…]
2:for $t=0$ to $H_{O}\times W_{O}+K-2$ do
3:  for $i=0$ to $K-1$ do
4:   for $j=0$ to $K-1$ do
5:    if $i=0$ then
6:     if $t=0$ then
7:      $PE_{i,j}(t)\leftarrow I_{ext}$
8:     else if $1\leq t\leq W_{O}-1$ then
9:      if $j=K-1$ then
10:       $PE_{i,j}(t)\leftarrow I_{ext}$
11:      else
12:       $PE_{i,j}(t)\leftarrow I_{R}$
13:      end if
14:     else if $t\bmod W_{O}=0$ then
15:      $\alpha\leftarrow\alpha+1$
16:      $PE_{i,j}(t)\leftarrow I_{D}$
17:     else if $t=\alpha W_{O}+1$ then
18:      if $0\leq j<K-1$ then
19:       $PE_{i,j}(t)\leftarrow I_{R}$
20:      else
21:       $PE_{i,j}(t)\leftarrow I_{D}$
22:      end if
23:     else if $t\geq H_{O}\times W_{O}$ then
24:      $PE_{i,j}(t)\leftarrow Idle$
25:     else
26:      if $0\leq j<K-1$ then
27:       $PE_{i,j}(t)\leftarrow I_{R}$
28:      else
29:       if $W_{I}\leq 2K$ then
30:        $PE_{i,j}(t)\leftarrow I_{ext}$
31:       else
32:        if $(\alpha+1)W_{O}-K<t\leq(\alpha+1)W_{O}-1$ then
33:         $PE_{i,j}(t)\leftarrow I_{ext}$
34:        else
35:         $PE_{i,j}(t)\leftarrow I_{D}$
36:        end if
37:       end if
38:      end if
39:     end if
40:    else if $i=K-1$ then
41:     if $t<K-1$ then
42:      $PE_{i,j}(t)\leftarrow Idle$
43:     else if $t=K-1$ or $(t-K+1)\bmod W_{O}=0$ then
44:      $PE_{i,j}(t)\leftarrow I_{ext}$
45:     else
46:      if $0\leq j<K-1$ then
47:       $PE_{i,j}(t)\leftarrow I_{R}$
48:      else
49:       $PE_{i,j}(t)\leftarrow I_{ext}$
50:      end if
51:     end if
52:    else
53:     $PE_{i,j}(t)\leftarrow PE_{0,j}(t-i)$, with $t\geq i$
54:    end if
55:   end for
56:  end for
57:end for
Psums are accumulated vertically, thus from each $PE_{i,j}$ to each $PE_{i+1,j}$. Eventually, an adder tree accumulates the psums coming from the bottom $PE_{K-1,j}$ elements. For very small ifmaps, with $W_{I}\leq 2K$, the number of registers may be lower than $K$, thus the diagonal connections may also interest some of the PEs.
@@
To better understand how TrIM works, we consider the case reported in Fig. 5, where a $\times 5$ 5 ifmaps, with $I=1,2,...,25$ being the inputs, and a $\times 3$ 3 kernel, with $W=A,B,...,I$ being the weights, are considered. The equivalent TrIM architecture is made of 3 rows (named $Row_{0}$, $Row_{1}$, and $Row_{2}$), each having 3 PEs. The leftmost PEs belonging to $Row_{1}$ and $Row_{2}$ are connected to SRBs with depth 1. A final adder tree, consisting of two adders, manages the psums coming from the PEs of $Row_{2}$. Preliminarily, 3 cycles are needed to load the $\times 3$ 3 weights, arranged in rows of 333 weights per cycle. Once all the weights are stored, all the convolutions are performed, as detailed in Fig. 5. For each cycle, the numbers in the PEs indicate the current computation (new multiplications in bold). The different colors refer to the input movement that has led to that multiplication: cells in blue refer to vertical input movement; cells in red refer to right-to-left movements; cells in orange refer to diagonal movements. Input activations temporarily stored in SRBs are reported using white boxes. Finally, the green boxes report the final output from the adder tree. To facilitate the reading of Fig. 5, we pick the case of cycle 4 as an example: $Row_{0}$ multiplies $I=6,7,8$ with $W=A,B,C$, where $I=6$ is reused diagonally from $SRB_{0}$, while $I=7,8$ are reused diagonally from $PE_{1,0}$ and $PE_{1,1}$, respectively. $Row_{1}$ multiplies the inputs $I=8,9,10$ with $W=D,E,F$, and accumulates these with the psums coming from $Row_{0}$. While $I=8,9$ are reused through right-to-left movements, $I=10$ is supplied externally. $Row_{2}$ multiplies the inputs $I=12,13,14$ with $W=G,H,I$, and accumulates these with the psums coming from $Row_{1}$. While $I=12,13$ are reused through right-to-left movements, $I=14$ is supplied externally. At this cycle, $SRB_{0}$ and $SRB_{1}$ receive $I=7$ and $I=11$, respectively. Finally, the psums coming from $PE_{2,0}$, $PE_{2,1}$, and $PE_{2,2}$ are accumulated through the adder ($O_{0,0}$ result).
@@
Overall, the case reported in Fig. 5 allows us to appreciate how TrIM maximizes input utilization. In fact, the number of total memory accesses from the main memory is 29, of which only 4 accesses refer to inputs read more than once. To better spotlight this, let consider the case $I=13$. This input is read once from the main memory (cycle 3) and reused 8 times (from cycle 4 to cycle 9) through right-to-left and diagonal movements.
III-B Support for Fully-Connected Layers
The TrIM-based SA is tailored for convolution workloads, overcoming the data redundancy issue introduced by Conv-to-GeMM. However, the architecture can be adapted to support conventional FCLs. In this case ifmaps are treated as one-dimensional vectors. In such a scenario, TrIM behaves as a WS-based SA: weights are loaded in the preliminary stage (from top to bottom, through vertical links). Then, new inputs are supplied to the PEs through parallel vertical injection. Horizontal and diagonal movements are not exploited. Psums are vertically accumulated and later reduced through the adder tree at the bottom of the SA. Overall, data are processed in a pipeline fashion without interruptions over time. During FCLs’ execution, SRBs are not exploited. However, considering that CNNs exploit hundreds of CLs and few FCLs[1], the overhead of SRBs in FCL is not a disadvantage in general.
III-C Differences with WS and RS Dataflows
@@
  • Data redundancy: WS has data redundancy at the main memory level, which adds a significant energy overhead. RS has data redundancy at the array level through memory blocks, adding still significant energy overhead. While RS moves the redundancy penalty from the off-chip to the on-chip level, TrIM overcomes the input redundancy problem and, moreover, maximizes the data utilization by recycling the same input as much as possible.
@@
  • Array sizes: WS and TrIM make use of $K\times K$ PEs. In contrast, RS is a rectangular array with sizes $K\times H_{O}$, thus exhibiting more limitations regarding the design scalability and area efficiency. In addition, the presence of scratch pads at the PE level makes the RS even bigger.
@@ IV An Analytical Model for Systolic Arrays
In order to highlight the advantages provided by TrIM, we built an analytical model to explore and evaluate the design space of TrIM, searching for the optimum design points that lead to the physical implementation phase. The aim of this model is twofold: (i) providing insights about memory accesses, throughput, and local buffers/registers; (ii) comparing TrIM to WS and RS dataflows. Without loss of generality, all the metrics refer to a convolution between a $H_{I}\times W_{I}$ ifmap and a $K\times K$ kernel. However, these can be easily scaled to manage $M$ ifmaps and $N$ filters of $M$ kernels.
@@ IV-B Modelling TrIM Dataflow
[…]
@@
The component $4\times K^{2}$ relates to the registers available at the PE level (4 for each PE). The second block $(K-1)\times(W_{I}-K-1)$ refers to the registers that model the SRBs. Finally, the extra register equips the adder tree.